How long does it take for a deposit of $ 900 to double at 4​% compounded​ continuously?
Answer by solve_for_x(190)

The doubling time can be found from:

t = ln(2) / rate

with the rate expressed as a decimal.

With a rate of 4% = 0.04, the doubling time would be:

t = ln(2) / 0.4 = 17.32 years